# Useful entanglement can be extracted from noisy graph states

### v1

@@ 3 Correlations and measurement
The remaining qubits are fully defined by the correlations developed during the measurement process. Thus, it is possible to find measurement schemes for which the output state has certain properties: e.g., if the goal is to extract a Bell pair from a larger graph state |$\ket{G}$, a measurement scheme $\prod_{i\in I}o_{i}$ must be found such that there exist stabilizers $Q$ of |$\ket{G}$ such that $Q_{I\cup J}=O_{I\cup J}$ and ensuring Bell-like correlations between the terminal vertices. For example in the 5 qubit line graph and the above chosen measurements, the Bell state stabilizers are extracted; more elaborate measurement patterns have been studied for use in quantum networks [20, 19, 27]. Further examples for graph families analyzed later in this article can be found in Appendix A.
@@
If a stabilizer operator $Q=\pm\prod_{j\in J}q_{j}$ is fully embedded in the measurement scheme $O=\prod_{i\in I}o_{i}$ such that $J\subset I$ and $Q=\pm O|_{J}$, the only observable sign structures within $J$ are those such that $\prod_{j\in J}s_{j}=\pm 1$, where the sign is determined by the sign in $Q$.
@@
As mentioned in the introduction, the creation of a Bell pair is equivalent to the task of information transfer (teleportation) across the graph. This is so because if the Bell pair creation in the graph $G$ is viewed in the context of a larger graph $G^{\prime}$, this process (followed by measurement of one of the terminal qubits) performs qubit fusion: the unmeasured terminal qubit acts as the two combined. Since all local measurements with nonoverlapping supports do commute, this can be done before any other measurement or afterwards: in the latter case, the information created in one of the qubits is merged with the second one. This is captured by the following Lemma, represented pictorially by the Figure 3.
Lemma 2.
Consider a graph $G=\{V,E\}$, and choose the two terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ such that $V=\{{\operatorname{in}},{\operatorname{out}}\}\cup I$. Suppose that the corresponding graph state |$\ket{G}$ undergoes a sequential measurement of $I$ producing a Bell pair across $\{{\operatorname{in}},{\operatorname{out}}\}$ in the meaning of Lemma 1.
With this determined, let us further assume that $G^{\prime}=(V^{\prime},E^{\prime})$ embeds $G$ in such a way that $\{{\operatorname{in}},{\operatorname{out}}\}$ is a separating set: the only connections between $G$ and the rest of $G^{\prime}$ are at these two vertices. Let us denote the vertices in $V^{\prime}\setminus V$ to which both qubits are connected by $\mathcal{N}^{\prime}({\operatorname{in}})$ and $\mathcal{N}^{\prime}({\operatorname{out}})$, respectively, and assume that the two sets are disjoint.
Within $G^{\prime}$, the aforementioned Bell pair-generating measurement followed by a product of single qubit unitary operations on the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ and the measurement $X_{\operatorname{in}}$ effectively performs qubit fusion of the terminal qubits. The new state is a graph state where the qubits $V^{\prime}\setminus V$ are intact, and the new neighborhood of the remaining qubit is $\mathcal{N}_{{\operatorname{out}}}\cup\mathcal{N}_{{\operatorname{in}}}$
Proof.
If within $G$ a Bell pair is produced across $\{{\operatorname{in}},{\operatorname{out}}\}$ as a result of the measurement pattern $O=\prod_{i\in L}o_{i}$ (where $L=V\setminus\{{\operatorname{in}},{\operatorname{out}}\}$), this means that there exist three $O$-consistent stabilizers involving the terminal vertices in the sense of Lemma 1. Combinatorical considerations show that one of those must be of a form such that the operators at the terminal vertices are $A_{\operatorname{in}},B_{\operatorname{out}}\in\{X,Y\}$ (but not $Z$), chosen independently.
Thus, there exists an $G$-stabilizer operator of the form ±$\pm A_{\operatorname{in}}B_{\operatorname{out}}\prod_{i\in I}o_{i}=g_{\operatorname{in}}g_{\operatorname{out}}\prod_{i\in I^{\prime}}g_{i}$, with $I,I^{\prime}\subset L$. Note that the definition of $g_{i}$ depends on the graph in question (Equation 6): if by $g^{\prime}_{i}$ we denote the generators corresponding to the entire graph $G^{\prime}$, the result is also a stabilizer, differing only by the $Z$ operators in the set $\mathcal{N}^{\prime}({\operatorname{in}})\cup\mathcal{N}^{\prime}({\operatorname{out}})$.
After the $A_{\operatorname{in}}$-basis measurement, we get a new stabilizer provided by Lemma 1:
(±$\left(\pm s_{\operatorname{in}}\prod_{i\in L}s_{i}\right)B_{\operatorname{out}}\mathchoice{}{}{}{}\prod_{j\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}\mathchoice{}{}{}{}Z_{j},$ (14)
which can be taken to be a new stabilizer generator associated with the vertex ${\operatorname{out}}$, and brought to the canonical form of $X_{\operatorname{out}}\prod_{j\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}Z_{j}$ by local unitary basis change of the ${\operatorname{out}}$-qubit. The properties of other stabilizers involving the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ can be proven similarly.
Thus, after the measurement within the Bell-generating part $L\cup\{{\operatorname{in}}\}$, the ${\operatorname{out}}$-qubit behaves exactly like it was connected to the neighborhood $\mathcal{N}^{\prime}_{\operatorname{in}}$ as well: the two vertices are fused and the information between them is transferred. ∎
@@ 4 Imperfect graph states
The reasoning presented above assumes that perfect $\operatorname{CZ}$ gates can be performed. However, this is not the case in experimental setups: some kind of noise is always present. In this section, we analyze three classes of physically motivated noise models; they all stem from different methods of graph state preparation. The origins and effects are sketched; for the detailed derivation, please refer to Appendix B.
Regardless of origin, noise processes often have a stabilizer-consistent description: vaguely speaking, they produce a mixture of stabilizer states, or the noise effect can be otherwise understood using the stabilizer-like correlations. In later sections, we show how to use this observation to mitigate the effects of noise.
@@ 4.1 Uncorrelated edge noise
Implementations like superconducting qubit quantum computers [28, 29], or ion traps [30, 31] prepare graph states by letting an initial product state evolve via an engineered Ising-like interaction Hamiltonian:
$H=\sum_{i\neq j}\gamma_{i,j}\frac{(Z_{i}+\operatorname{\mathds{1}})\otimes(Z_{j}+\operatorname{\mathds{1}})}{4}.$ (15)
Evolution according to such an interaction pattern effectively implements a sequence of controlled-phase operators $\operatorname{C\Phi}_{i,j}=\operatorname{diag}(1,1,1,\exp(-i\varphi_{i,j}))$. The structure of phases $\varphi_{i,j}=\gamma_{i,j}t$ is determined by the interaction strengths $\gamma_{i,j}$ and the evolution time $t$: inaccuracies in controlling either of them lead to the generation of a weighted graph state [22]. However, under general assumptions, the effective quantum channel implemented this way has a simple decomposition into local (one or two edges) operations, on top of the product of a sequence of $\operatorname{CZ}$ operations preparing the desired state.
Ideally, all of the generated phases $\varphi_{i,j}$ are equal to $\pi$. More realistically, with each experimental run, the actual realization will be $\varphi_{i,j}=\pi+\varepsilon_{i,j}$ with $\varepsilon_{i,j}\ll 1$. We discuss the two extremal cases of correlations of phase noises $\varepsilon_{i,j}$. The first, uncorrelated phase noise, assumes that the noises at different edges $(i,j)$ are completely independent. The other, correlated phase noise, assumes the contrary: each phase noise factor is equal to any other in a single experimental run, $\varepsilon_{i,j}=\varepsilon$. The noise encountered in experiments likely has yet a different structure with partial correlations between the edges; still, the two extremal processes may help in modeling the real-life scenario and strategies developed to mitigate them should apply in the more general cases.
Phases uncorrelated across the edges effectively lead to a probabilistic $\operatorname{CZ}$ (see the Appendix B) gate being implemented, if only each $\varepsilon_{i,j}$ is symmetrically distributed around 00 (so that $\varphi=\pi+\varepsilon$ is centered around $\pi$). Each of controlled-$Z$ unitary operators in Equation 1 is replaced with the quantum channel
@@
where $p$ depends on the distribution of $\varepsilon$ and $\operatorname{\mathcal{CZ}}[\rho]=\operatorname{CZ}\rho\operatorname{CZ}^{\dagger}$.
Thus effectively the effect of noise is the generation of randomized graph states [32]: ensemble of graph states built atop the original graph $G$ by removing the edge $(i,j)\in E$ with probability $p_{i,j}$; related graph ensembles have been introduced in [33] in a classical context. For the constant probability, the end state can be written as
$\rho=\sum_{E^{\prime}\subset E}(1-p)^{|E^{\prime}|}p^{|E\setminus E^{\prime}|}\ket{G^{\prime}}\bra{G^{\prime}},$ (17)
@@
Note that it is an effective description: in each experimental run, the prepared state is a pure weighted graph state [13]; the density operator in Equation 17 is a state of knowledge about the system, averaged over different noise realizations.
@@ 4.2 Correlated edge noise
If the weights in the Hamiltonian of Equation 15 are perfect ($\gamma_{i,j}=1$) but the interaction time is not perfectly controlled, within one experimental run, all the resulting phases $\varphi_{i,j}$ are equal. This is correlated phase noise: it has an effect similar, but more involved, to the uncorrelated case. Here, we also assume that $\varphi$ is symmetrically distributed around $\pi$. Let us assume that in each experimental run $\varphi_{i,j}=\varphi=\pi+\varepsilon$, and $\varepsilon$ is distributed with the normal distribution of the standard deviation $\sigma$. Then, with $p$ depending on the phase distribution (see Appendix B for details), the lowest order of approximation of the state (essentially, keeping only terms of order $\varepsilon^{2}$ in the end state and subsequent averaging) can be described with the help of square roots of the unitary operator $\operatorname{CZ}$:
$\rho=$ $(1-p\lvert E\rvert)\ket{G}\bra{G}+p\sum_{e\in E}\outerproduct{G-e}{G-e}$ | (18)
$+p\!\!\!\!\sum_{\{e,e^{\prime}\}\subset E\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \phantom{x}}^{e\neq e^{\prime}}\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\phantom{x}\leavevmode\nobreak\ \leavevmode\nobreak\ \pm_{1},\pm_{2}}\!\!\!\!\!\!\!\pm_{1}\!\pm_{2}\!{\operatorname{\mathcal{CS}}_{e}^{\pm_{1}}}\!\!{\operatorname{\mathcal{CS}}_{e^{\prime}}^{\pm_{2}}}[\ket{G}\bra{G}].$
Note that this decomposition is not convex and cannot be interpreted as an ensemble: the minus signs in front of the decomposition terms prevent this. Signs appear as a result of discrete decomposition of controlled-phase and subsequent averaging; see Appendix B for details. In this decomposition, $G-e\coloneqq(V,E\setminus\{e\})$ is the graph with edge removed, and ${\operatorname{\mathcal{CS}}^{\pm}}$ is the application of one of the two unitary square roots of $\operatorname{CZ}$:
where
$\sqrt{\operatorname{CZ}^{\pm}}\coloneqq\frac{1}{2}\left[\left(1+e^{\pm i\frac{\pi}{2}}\right)\operatorname{\mathds{1}}+\left(1-e^{\pm i\frac{\pi}{2}}\right)\operatorname{CZ}\right].$ (20)
@@ 4.3 Local $Z$ flip noise
In linear optic experiments, where Bell pairs are generated and subsequently fused to get a graph state, the entangling operation may fail as a result of partial photon distinguishability (by imperfect frequency of spatial mode overlap) [34, 35, 36] – for a derivation see the Appendix B. The effect of a noisy fusion gate described this way can be modeled as a perfect fusion followed by a probabilistic application of the $Z$ unitary to the surviving optical qubit $i$:
@@ 4.4 Quantification of the noise effects
In the following chapters, we determine the entanglement quality of a small (two- or three-qubit) subsystem left after a measurement procedure performed on a graph state: this is the Bell pair or GHZ state extraction, since ideally the resulting state would represent correlations found in these states. We have chosen fidelity with respect to the ideal state as a measure of the entanglement quality. In the absence of noise, the initial state |$\ket{G}$ is pure, and measurement of an $i$-th qubit in the basis of an operator $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ leads to observation of an outcome ±$\pm 1$. On the algebraic level, this corresponds to application of an operator $\frac{1}{2}\left(\mathbbm{1}\pm o_{i}\right)$ to the pre-measurement state, and after the entire procedure, the post-measurement state |$\ket{\psi}$ is still pure. This allows a simple determination of the fidelity with respect to the noisy state $\rho$: if the post-measurement state is $\sigma$, the following holds.
@@
Note that the ideal post-measurement state depends on the observed outcomes during the experiment: they do affect the signs in the final stabilizer structure (see Lemma 1). Therefore, the ideal state |$\ket{\psi}$ must be determined from the signs of the outcomes. This is only possible if the structure is consistent with the stabilizers of the ideal state; this is not an issue, since we explicitly postselect on the right set of outcomes in the following sections as part of the error mitigation strategy. Finally, the average fidelity $\langle\mathcal{F}\rangle$ is determined for all possible measurement outcomes consistent with the stabilizer structure; this is effectively approximated by taking $N$ samples from the defining ensembles for a specific noise model with a Monte Carlo algorithm and thus
$\langle\mathcal{F}\rangle=\frac{1}{N}\!\!\!\sum_{\text{outcomes}}\!\!\!\mathcal{F}(\text{outcome}).$ (23)
Since we are mostly interested in the behavior in the low noise limit, we define fidelity susceptibility $\alpha$ as the rate of change of mean fidelity $\langle\mathcal{F}\rangle(p)$: the ensemble of extracted and postselected states depends on the noise, and the susceptibility captures the effects of small but not negligible noise.
@@
For zero noise, the fidelity is equal to 1 by definition. For any other amount, the fidelity may only decrease, and hence the susceptibility is positive. Its numerical magnitude determines how fragile the extraction procedure is to the effects of noise: a robust one has a small (or, in some cases, even zero) susceptibility.
@@ 5 Extraction from noisy ensembles
In this section we apply the methods developed in previous parts to show the effects of the noise and find measurement patterns that minimize them. Instead of focusing on teleportation directly, we analyze the (equivalent) task of extracting a high-quality Bell pair from a larger noisy state, close to a graph state |$\ket{G}$.

### v2

@@ 3 Correlations and measurement
The remaining qubits are fully defined by the correlations developed during the measurement process. Thus, it is possible to find measurement schemes for which the output state has certain properties. For example, if the goal is to extract a Bell pair from a larger graph state |$\ket{G}$, a measurement scheme $\prod_{i\in I}o_{i}$ must be found such that there exist stabilizers $Q$ of |$\ket{G}$ such that $Q_{I\cap J}=O_{I\cap J}$ and ensuring Bell-like correlations between the terminal vertices. For example in the 4-qubit path graph of Figure 3a and the above chosen measurements, the Bell state stabilizers are extracted. More elaborate measurement patterns have been studied for use in quantum networks [12, 13, 40]. Further examples for graph families analyzed later in this article can be found in Appendix B.
@@
If a stabilizer operator $Q=\pm\prod_{j\in J}q_{j}$ is fully embedded in the measurement scheme $O=\prod_{i\in I}o_{i}$ such that $J\subset I$ and $Q=\pm O|_{J}$, the only observable sign structures within $J$ are those such that $\prod_{j\in J}s_{j}=\pm 1$, where the sign is determined by the sign $\pm$ in front of the product defining $Q$.
@@ 4 Imperfect graph states
In this section, we analyze three classes of physically motivated noise models. They all stem from different methods of graph state preparation. The origins and effects are only sketched here due to complexity, please refer to Appendix C for the detailed derivation.
As outlined in the previous section, for perfect graph states Bell pairs can be extracted, provided the vertices of interest are connected by a path. Once a path is found, vertices adjacent to it are measured in the $Z$-basis and then the path vertices in the $X$-basis. In Section 5 we show that in the presence of noise, the performance of a simple path graph is not optimal, and choice of other graphs can enhance entanglement quality of the remaining qubits.
As a minimal example, consider the graph shown in Figure 3b. It may arise during preparation of path graph (Figure 3a) via an imperfect $\operatorname{CZ}$ gate, and in such case the Bell pair can not be generated: the two graph parts are uncorrelated, and local measurements can not produce entanglement from a product state. In a path graph, even a single missing edge leads to this scenario.
Regardless of origin, noise processes often have a stabilizer-consistent description. They produce a mixture of stabilizer states, or the noise effect can be otherwise understood using the stabilizer-like correlations. In Section 5, we show how to use this observation to mitigate the effects of noise.
Implementations like superconducting qubit quantum computers [41, 42], or ion traps [43, 44] prepare graph states by letting an initial product state evolve via an engineered Ising-like interaction Hamiltonian
$H=\sum_{i\neq j}\gamma_{i,j}\frac{(Z_{i}+\operatorname{\mathds{1}})\otimes(Z_{j}+\operatorname{\mathds{1}})}{4},$ (15)
where $\gamma_{i,j}$ is the interaction strength between qubits $i$ and $j$. Evolution according to such an interaction pattern effectively implements a sequence of controlled-phase operators $\operatorname{C\Phi}_{i,j}=\operatorname{diag}(1,1,1,\exp(-i\varphi_{i,j}))$. The structure of phases $\varphi_{i,j}=\gamma_{i,j}t$ is determined by the interaction strengths $\gamma_{i,j}$ and the evolution time $t$. Inaccuracies in controlling either of them lead to the generation of a weighted graph state [34]. However, under general assumptions, the effective quantum channel implemented this way has a simple decomposition into local (one or two edges) operations, on top of the product of a sequence of $\operatorname{CZ}$ operations preparing the desired state.
Ideally, all of the generated phases $\varphi_{i,j}$ are equal to $\pi$. More realistically, with each experimental run, the actual realization will be $\varphi_{i,j}=\pi+\varepsilon_{i,j}$ with $\varepsilon_{i,j}\ll 1$. We discuss the two extremal cases of correlations of phase noises $\varepsilon_{i,j}$. The first, uncorrelated phase noise, assumes that the noises at different edges $\{i,j\}$ are completely independent. The other, correlated phase noise, assumes the contrary: each phase noise factor is equal to any other in a single experimental run, $\varepsilon_{i,j}=\varepsilon$. The noise encountered in experiments likely has yet a different structure with partial correlations between the edges; still, the two extremal processes may help in modeling the real-life scenario and strategies developed to mitigate them should apply in the more general cases.
@@ 4.1 Uncorrelated edge noise
The effect of phase noise is especially simple if the phases $\varphi$ are distributed symmetrically and uncorrelated around $\pi$. The resulting channel is described by a probabilistic application of $\operatorname{CZ}$: with probability $(1-p)$, an edge is created via $\operatorname{CZ}$, and with probability $p$, no operation is performed. Each of controlled-$Z$ unitary operators in Equation 1 is replaced with the quantum channel
@@
[…]
$\rho=\sum_{E^{\prime}\subset E}(1-p)^{|E^{\prime}|}p^{|E\setminus E^{\prime}|}\outerproduct{G^{\prime}}{G^{\prime}},$ (17)
@@
Note that it is an effective description: in each experimental run, the prepared state is a pure weighted graph state [21]. The density operator in Equation 17 is a state of knowledge about the system, averaged over different noise realizations.
@@ 4.2 Correlated edge noise
If the weights in the Hamiltonian of Equation 15 are all equal ($\gamma_{i,j}=1$) but the interaction time is not perfectly controlled, within one experimental run, all the resulting phases $\varphi_{i,j}$ are equal. This is the correlated phase noise: it has an effect similar, but more involved, to the uncorrelated case. Here, we also assume that $\varphi$ is symmetrically distributed around $\pi$. Let us assume that in each experimental run $\varphi_{i,j}=\varphi=\pi+\varepsilon$, and $\varepsilon$ is distributed with the normal distribution of the standard deviation $\sigma$. The resulting state can be described by a parameter $p$ depending on the distribution of $\varepsilon$ (Equation 34). By keeping only the lowest nontrivial order of noise effects and subsequent averaging, the end state can be described with the help of square roots of the unitary operator $\operatorname{CZ}$:
[…]
$+\frac{p}{2}\sum_{\begin{subarray}{c}e,e^{\prime}\in E\\ e\neq e^{\prime}\end{subarray}}\sum_{s,s^{\prime}\in\{\pm 1\}}ss^{\prime}\operatorname{\mathcal{CS}}_{e}^{s}\operatorname{\mathcal{CS}}_{e^{\prime}}^{s^{\prime}}[\outerproduct{G}{G}],$
where the summation runs over graphs $G^{\prime}\subseteq G$ with single edges removed from $G$, and ${\operatorname{\mathcal{CS}}^{\pm}}[\rho]$ is the application of one of the two unitary square roots of $\operatorname{CZ}$. See Appendix C for details.
@@ 4.3 Local $Z$ flip noise
In linear optic experiments, where Bell pairs are generated and subsequently fused to get a graph state, the entangling operation may fail as a result of partial photon distinguishability This can be caused by imperfect frequency or spatial mode overlap [47, 48, 49]. See LABEL:{app:noise} for a derivation. The effect of a noisy fusion gate described this way can be modeled as a perfect fusion followed by a probabilistic application of the $Z$ unitary to the surviving optical qubit $i$:
@@ 4.4 Quantification of the noise effects
In the following sections, we determine the entanglement quality of two- and three-qubit Bell and GHZ graph states left after a measurement procedure performed on a graph state. To quantify the entanglement quality of the post-measurment state, we use the fidelity with respect to the ideal state. If the post-measurement state is $\sigma$, the following holds:
@@
Note that the ideal post-measurement state depends on the observed outcomes during the experiment: they do affect the signs in the final stabilizer structure (see Lemma 1). Therefore, the ideal state |$\ket{\psi}$ must be determined from the signs of the outcomes, provided they are consistent with the stabilizer structure. The different outcomes may appear with varying probabilities, and the expectation value of associated fidelities (Equation 20) is the average fidelity $\langle\mathcal{F}\rangle$. This is effectively approximated by taking $N$ samples from the defining ensembles for a specific noise model with a Monte Carlo algorithm and thus
$\langle\mathcal{F}\rangle\approx\frac{1}{N}\!\!\!\sum_{\text{outcomes}}\!\!\!\mathcal{F}(\text{outcome}).$ (21)
Since we are mostly interested in the behavior in the low noise limit, we define fidelity susceptibility $\alpha$ as the rate of change of mean fidelity $\langle\mathcal{F}\rangle(p)$: the ensemble of extracted and postselected states depends on the noise, and the susceptibility captures the effects of small but not negligible noise:
@@
For zero noise, the fidelity is equal to 1 by definition. For any other amount, the fidelity may only decrease, and hence the susceptibility is positive. Its numerical magnitude determines how fragile the extraction procedure is to the effects of noise: a robust one has a small susceptibility.
Practically, the fidelity susceptibility is estimated here by a discrete derivative approximation. The fidelity $\langle\mathcal{F}\rangle_{*}$ is determined by choosing a small value of noise $p_{*}$ and computing the fidelity for it.
Finally, the fidelity susceptibility is approximated as $\alpha\approx\frac{1-\langle\mathcal{F}\rangle_{*}}{p_{*}}$. In our computations, we chose $p_{*}\approx 10^{-2}$. The noise value $p_{*}$ were chosen such that it is empirically low enough for reasonable approximation and high enough such that numerical noise is not relevant.
@@ 5 Extraction from noisy ensembles
In this section we apply the methods developed in previous parts to show the effects of the noise and find measurement patterns that minimize them. As mentioned previously, we analyze task related to teleportation: extraction of a high quality Bell pair from a larger noisy state, close to a graph state |$\ket{G}$. Various graph families are introduced together with appropriate measurement patterns, and the effect of different types of noise is analyzed for all of them. The main results are shown in Figures 6 and 7.
Monte Carlo sampling was used to calculate the numerical results presented in this section. We mostly concentrate on the scenario where the entangling gates are almost perfect, and in such case the fidelity susceptibility (Equation 22) quantifies the noise effects. Practically, the average fidelity (Equation 21) is determined by simulating the random outcomes of quantum measurement (with postselection where appropriate), and so is the fidelity susceptibility $\alpha$. For the estimates of $\alpha$ only a limited number of quantum states forming the ensemble (discussed in Section 4) is relevant, and for calculation only this subset is taken into account, reducing the uncertainty inherent to Monte Carlo procedure. The code used for calculation is publicly available [32], the numerical details can be found there.